Suppose there is an isolated charge (+1.5π‘₯10βˆ’19𝐢) that is displaced from point A to B due to the electric force. As the charge moves, it produces an electric field which has a magnitude of 6.9 N/C. Find the work done by the electric force of the charge if it moves 1.5 m from its original position.


1
Expert's answer
2022-03-04T08:31:03-0500

The net work done

"W=qEd\\
=1.5*10^{-19}*6.9*1.5=1.55*10^{-18}\:\rm J"